# Geometry of convex geometries

### v1

@@ 3.1 Realizability of COMs
Let $(\mathcal{H}^{\prime},K^{\prime})$ be a realization of $\mathcal{M}$ in ${\mathbb{R}}^{d}$ as in Lemma 3. Since the vectors normal to the central hyperplanes of $\mathcal{H}^{\prime}$ are independent, they form a basis of ${\mathbb{R}}^{d}$, thus there exists a linear transformation $f$ of ${\mathbb{R}}^{d}$ mapping these vectors to the coordinate vectors. Then $f$ maps the hyperplanes of $\mathcal{H}^{\prime}$ to the coordinate hyperplanes $\mathcal{H}_{d}$ and the open polyhedron $K^{\prime}$ to an open polyhedron $K^{\prime\prime}$. Furthermore, each open cell $C^{\prime}$ of $K^{\prime}$ bounded by a set $\mathcal{H}_{0}^{\prime}$ of hyperplanes of $\mathcal{H}^{\prime}$ is mapped to a nonempty open cell $C^{\prime\prime}$ of $K^{\prime\prime}$ bounded by the set $\mathcal{H}_{0}^{\prime\prime}$ of hyperplanes of $\mathcal{H}_{d}$, which are images of the hyperplanes of $\mathcal{H}_{0}^{\prime}$. Consequently, $(\mathcal{H}_{d},K^{\prime\prime})$ realizes $\mathcal{M}$. ∎
@@ 3.2 General properties of realizability
A hyperplane arrangement of $d$ hyperplanes in $\mathbb{R}^{d-1}$ has less than $2^{d}$ maximal cells, see e.g., [5, Exercise 4.3.5]. Thus, the free COM (which has $2^{d}$ topes) cannot be realized in $\mathbb{R}^{d-1}$. ∎
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be an OM and suppose that it can be realized as a (realizable) COM $\mathcal{M}=\mathcal{M}(\mathcal{H},K)$ in $\mathbb{R}^{d}$. Suppose without loss of generality that $K$ is full-dimensional, otherwise we project into the affine hull of $K$. Since $\mathcal{M}$ is an OM, we have that $\mathbf{0}\in\mathcal{L}$ (recall that $\mathbf{0}$ is the all-zeros vector). Hence, there is a point $x\in K$ in which all the hyperplanes of $\mathcal{H}$ intersect. After translating $x$ to the origin of $\mathbb{R}^{d}$ we see that $\mathcal{H}$ is a central hyperplane arrangement. Hence, its combinatorics is determined by any arbitrary small sphere $S$ around $x$, see e.g. [5, Section 1.2.(c)]. Therefore, we can set $K=\mathbb{R}^{d}$ and get a a representation of $\mathcal{M}$ by a central hyperplane arrangement. ∎
@@ 3.3 Generalized convex shellings
Convex geometries are considered as generalizations of Euclidean convexity in the following sense. Let $\mathcal{A}=(U,\mathcal{C})$ be a convex geometry. We say that $\mathcal{A}$ is CG-realizable if $U$ can be viewed as a finite subset of points of ${\mathbb{R}}^{d}$ such that $X\subseteq U$ belongs to $\mathcal{C}$ if and only if the Euclidean convex hull of $X$ does not contain other points of $U$: $\operatorname{conv}_{{\mathbb{R}}^{d}}(X)\cap U=X$. Not every convex geometry can be realized in this way. Kashiwabara et al. [16] gave a realization theorem for all convex geometries using the notion of generalized convex shellings. Let $P$ and $Q$ be two finite sets of ${\mathbb{R}}^{d}$ such that the Euclidean convex hull of $Q$ does not intersect $P$. Then the family of subsets $\mathcal{C}(P,Q)=\{X\subseteq P:\operatorname{conv}(X\cup Q)\cap P=X\}$ of $P$ is called the generalized convex shelling on $P$ with respect to $Q$. It is easy to see that any generalized convex shelling is a convex geometry. Conversely, Kashiwabara et al. [16] proved that any convex geometry arises as a generalized convex shelling. If $Q=\varnothing$, then this leads to the notion of CG-realizability. If ${\mathcal{C}}=\mathcal{C}(P,Q)$, then we will say that $\mathcal{C}(P,Q)$ is a generalized convex shelling of $\mathcal{A}=(P,\mathcal{C})$. Richter and Rogers [24] proved that any convex geometry of (combinatorial) dimension $d$ has a generalized convex shelling in ${\mathbb{R}}^{d}$.
@@
Trivially, $P\in\mathcal{M}(P,Q)$. We say that $(P,Q)$ is a point representation of a COM $\mathcal{M}$ if $\mathcal{M}=\mathcal{M}(P,Q)$, i.e., the set of topes of $\mathcal{M}$ consists precisely of the $\{-1,+1\}$-vectors of all sets of $\mathcal{M}(P,Q)$. Now, we establish a correspondence between (hyperplane) realizations and point realizations of COMs and a link with generalized convex shellings. For this purpose we define a COM to be acyclic if it has the all plus-vector $(+1,\ldots,+1)$ as a tope.
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be an acyclic realizable COM. We have $\mathcal{M}=\mathcal{M}(\mathcal{H}^{\prime},K^{\prime})$ where ${\mathcal{H}}^{\prime}$ is a central hyperplane arrangement ${\mathcal{H}}^{\prime}$ and $K^{\prime}$ an open polyhedral cone in $\mathbb{R}^{d+1}$ as in Lemma 3. Now, the vectors $v_{e}$, normal to the hyperplanes $H^{\prime}_{e},e\in U$ of the arrangement ${\mathcal{H}}^{\prime}$, gives raise to a set ${\mathcal{V}}$ of vectors of ${\mathbb{R}}^{d+1}$. Analogously, the vectors $u_{j}$, normal to the hyperplanes $S^{\prime}_{j}$ of ${\mathcal{S}}^{\prime}$ defining the polyhedron $K^{\prime}$ and such that their product with the vectors with ends inside $K^{\prime}$ is positive, gives raise to a set ${\mathcal{U}}$ of vectors of ${\mathbb{R}}^{d+1}$. For a cell $C$ of the initial realization, each point $p$ of $C$ defines a vector $v_{p}$ belonging to the cone $C^{\prime}$. Let $H_{v_{p}}$ be the central hyperplane of ${\mathbb{R}}^{d+1}$ having $v_{p}$ as the normal vector. Then all vectors of ${\mathcal{U}}$ belong to a halfspace defined by $H_{v_{p}}$. Analogously, each vector $v_{e}\in{\mathcal{V}}$ belongs to the positive side of the hyperplane $H_{v_{p}}$ if and only if the cell $C$ is located on the positive side of the hyperplane $H_{e}\in{\mathcal{H}}$ (and thus the cone $C^{\prime}$ is on the positive side of the central hyperplane $H^{\prime}_{e}\in{\mathcal{H}}$). Therefore, the topes of $\mathcal{M}$ can be identified (in the set-theoretical language) with the subsets $X$ of $U$ such that there exists a central hyperplane $H^{\prime}$ of ${\mathbb{R}}^{d+1}$ such that all $U$ belong to one halfspace defined by $H^{\prime}$ and a vector $v_{e},e\in U$ belongs to the same halfspace defined by $H^{\prime}$ if and only if $e\in X$.
@@
Therefore, the hyperplane realizations and point representations of COMs are much simpler than generalized convex shellings and correspond to classical OM-realizations. The main results of our paper (Theorems 25 and 34) must be considered from this point of view, while compared to the results of [16] and [24] in the case of convex geometries. On the other hand, the generalized convex shellings characterize convex geometries [16].
A relaxed convex shelling is defined as a generalized convex shelling without requiring that $\operatorname{conv}(Q)\cap P=\varnothing$. Analogously to generalized convex shellings, relaxed convex shellings can be reformulated as follows: given two arbitrary finite sets $P$ and $Q$ of ${\mathbb{R}}^{d}$, a set $X$ belongs to $\mathcal{C}(P,Q)=\{X\subseteq P:\operatorname{conv}(X\cup Q)\cap P=X\}$ if and only if there exists of a set of halfspaces $H^{+}_{1},\ldots,H^{+}_{k}$, all containing $Q$ and such that $X=P\cap(\cap H_{i}^{+})$. A COM $\mathcal{M}$ is called intersection-closed if the set of topes of $\mathcal{M}$ defines an intersection-closed family of sets and $\varnothing\in\mathcal{M}$. Obviously, convex geometries are acyclic and their ideals are intersection closed COMs. Note however, that acyclicity corresponds to the universe being convex.
Lemma 12.
Any point representation $(P,Q)$ of an acyclic intersection-closed COM $\mathcal{M}$ yields a relaxed convex shelling representation of $\mathcal{M}$. If $\mathcal{M}$ is a convex geometry, then a relaxed convex shelling is a generalized convex shelling.
@@
Since $(P,Q)$ is a point representation of $\mathcal{M}$, we have $\mathcal{M}=\mathcal{M}(P,Q)$. From the definition of the point representations and relaxed convex shellings, we conclude that $\mathcal{M}(P,Q)\subseteq\mathcal{C}(P,Q)$. By Remark 11, $\mathcal{C}(P,Q)$ is closed by taking intersections. To prove the converse inclusion $\mathcal{C}(P,Q)\subseteq\mathcal{M}$, pick any set $X$ from $\mathcal{C}(P,Q)$. Then there exists a set $H^{+}_{1},\ldots,H^{+}_{k}$ of halfspaces, all containing $Q$ and such that $X=P\cap(\cap H_{i}^{+})$. Let $X_{i}=P\cap H_{i}^{+},i=1,\ldots,k$. Then each $X_{i}$ belongs to $\mathcal{M}(P,Q)$. Since $\mathcal{M}(P,Q)=\mathcal{M}$, each $X_{i}$ belongs to $\mathcal{M}$. Since $\mathcal{M}$ is intersection-closed and $\cap_{i=1}^{k}X_{i}=X$, we conclude that $X\in\mathcal{M}$, establishing that $\mathcal{C}(P,Q)\subseteq\mathcal{M}$. Since $\varnothing\in\mathcal{M}$, necessarily $\operatorname{conv}(Q)\cap P=\varnothing$, establishing that the relaxed convex shelling is a generalized convex shelling. ∎
@@ 4.1 Rooted circuits
Korte and Lovász [20] (see also the book [21]) established that antimatroids can be reconstructed from a unique minimal subset of rooted circuits, which they call critical rooted circuits. In terms of convex geometries, a rooted circuit $(C,a)$ of a convex geometry $\mathcal{A}=(U,\mathcal{C})$ is a critical rooted circuit if $\operatorname{conv}(C)\setminus\{a\}$ does not belong to $\mathcal{C}$ but $\operatorname{conv}(C)\setminus\{a,b\}$ belongs to $\mathcal{C}$ for any $b\in C\setminus\{a\}$. Denote by $\mathcal{R}_{0}(\mathcal{A})$ the set of all critical rooted circuits of $\mathcal{A}$.
Theorem 15 ([20, 21]).
Let $\mathcal{A}=(U,\mathcal{C})$ be a convex geometry. Then $\mathcal{A}$ can be reconstructed from the family $\mathcal{R}_{0}(\mathcal{A})$ of its critical rooted circuits. Furthermore, if $\mathcal{A}$ can be reconstructed from $\mathcal{R}^{\prime}\subseteq\mathcal{R}(\mathcal{A})$, then $\mathcal{R}_{0}(\mathcal{A})\subseteq\mathcal{R}^{\prime}$.
@@ 4.2 Positive circuits
Let $\mathcal{C}^{\prime\prime}=\{X\in\mathcal{C}:\forall P\in\mathcal{P}(\mathcal{I}),P\not\subseteq X\}$. Observe that $\mathcal{C}^{\prime\prime}$ is trivially included in $\mathcal{C}$. For any $X\in\mathcal{C}^{\prime}$ and any $P\subseteq U$, if $P\subseteq X$, then $P\notin\mathcal{P}(\mathcal{I})$. Consequently, $X\in\mathcal{C}^{\prime\prime}$ and thus $\mathcal{C}^{\prime}\subseteq\mathcal{C}^{\prime\prime}$.
Suppose now that there exists $Y\in\mathcal{C}^{\prime\prime}\setminus\mathcal{C}^{\prime}$ and assume that $Y$ is such a set of minimal size. By minimality of $Y$, for any $e\in Y$, $Y\setminus\{e\}\in\mathcal{C}$ if and only if $Y\setminus\{e\}\in\mathcal{C}^{\prime}$. Let $P_{0}=\{e\in Y:Y\setminus\{e\}\in\mathcal{C}^{\prime}\}$. Since $\mathcal{C}^{\prime}$ is intersection-closed, we conclude that for any $\varnothing\subsetneq P^{\prime}\subseteq P_{0}$, we have $Y\setminus P^{\prime}\in\mathcal{C}^{\prime}$. Consequently, $2^{P_{0}}\setminus\{P_{0}\}\subseteq\mathcal{C}^{\prime}_{|P_{0}}$. Since $Y\cap P_{0}=P_{0}$ and $Y\in\mathcal{C}^{\prime\prime}$, necessarily $P_{0}\notin\mathcal{P}(\mathcal{I})$. Consequently, there exists $X\in\mathcal{C}^{\prime}$ such that $P_{0}\subseteq X$. Let $Z=X\cap Y$ and observe that $P_{0}\subseteq Z$. Since $\mathcal{C}$ is intersection-closed, $Z\in\mathcal{C}$. If $Z=Y$, then $Y\subseteq X\in\mathcal{C}^{\prime}$, contradicting the fact that $\mathcal{C}^{\prime}$ is an ideal of $\mathcal{C}$. Thus, by the minimality of $Y$, we have $Z\subsetneq Y$. Since $Z\in\mathcal{C}^{\prime}$ and $Y\notin\mathcal{C}^{\prime}$, there exists a set $Z\subseteq A\subseteq Y$ and an element $e\in Y\setminus A$ such that $A\in\mathcal{C}^{\prime}$ and $B=A\cup\{e\}\in\mathcal{C}\setminus\mathcal{C}^{\prime}$ ($A$ and $e$ exist since $\mathcal{C}$ is ample and thus isometric). By minimality of $Y$, necessarily $B=Y$. Consequently $e\in P_{0}$, but this is impossible since $P_{0}\subseteq Z$. This establishes that $\mathcal{C}^{\prime\prime}\subseteq\mathcal{C}^{\prime}$. ∎

### v2

@@ 3.1 Realizability of COMs
Let (′′$(\mathcal{H}^{\prime\prime},K^{\prime\prime})$ be a realization of $\mathcal{M}$ in ${\mathbb{R}}^{d}$ provided by Lemma 5. The vectors normal to the central hyperplanes of $\mathcal{H}^{\prime\prime}$ generate ${\mathbb{R}}^{d}$. Now, we can extend $\mathcal{H}^{\prime\prime}$ to a hyperplane arrangement in a larger ${\mathbb{R}}^{n}$, so that the normal vectors of $\mathcal{H}^{\prime\prime\prime}$ are a basis of ${\mathbb{R}}^{n}$, while defining $K^{\prime\prime\prime}$ just as the restriction of $K^{\prime\prime}$ to ${\mathbb{R}}^{d}$. Now, there exists a linear transformation $f$ of ${\mathbb{R}}^{n}$ mapping these vectors to the coordinate vectors. Then $f$ maps the hyperplanes of $\mathcal{H}^{\prime\prime\prime}$ to the coordinate hyperplanes $\mathcal{H}_{n}$ and the relatively open polyhedron $K^{\prime\prime\prime}$ to a relatively open polyhedron $K$. Furthermore, each relatively open cell $C^{\prime\prime\prime}$ of $K^{\prime\prime\prime}$ bounded by a set $\mathcal{H}_{0}^{\prime\prime\prime}$ of hyperplanes of $\mathcal{H}^{\prime\prime\prime}$ is mapped to a nonempty cell $C$ of $K$ bounded by the set $\mathcal{H}_{0}$ of hyperplanes of $\mathcal{H}_{n}$, which are images of the hyperplanes of $\mathcal{H}_{0}^{\prime\prime\prime}$. Consequently, $(\mathcal{H}_{n},K)$ realizes $\mathcal{M}$. ∎
@@ 3.2 General properties of realizability
A hyperplane arrangement of $d$ hyperplanes in $\mathbb{R}^{d-1}$ has less than $2^{d}$ maximal cells, see e.g., [7, Exercise 4.3.5]. Thus, the free COM (which has $2^{d}$ topes) cannot be realized in $\mathbb{R}^{d-1}$. ∎
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be an OM and suppose that it can be realized as a (realizable) COM $\mathcal{M}=\mathcal{M}(\mathcal{H},K)$ in $\mathbb{R}^{d}$. Suppose without loss of generality that $K$ is full-dimensional, otherwise we project into the affine hull of $K$. Since $\mathcal{M}$ is an OM, we have that $\mathbf{0}\in\mathcal{L}$ (recall that $\mathbf{0}$ is the all-zeros vector). Hence, there is a point $x\in K$ in which all the hyperplanes of $\mathcal{H}$ intersect. After translating $x$ to the origin of $\mathbb{R}^{d}$ we see that $\mathcal{H}$ is a central hyperplane arrangement. Hence, its combinatorics is determined by any arbitrary small sphere $S$ around $x$, see e.g. [7, Section 1.2.(c)]. Therefore, we can set $K=\mathbb{R}^{d}$ and get a a representation of $\mathcal{M}$ by a central hyperplane arrangement. ∎
@@ 3.3 Generalized convex shellings
Convex geometries are considered as generalizations of Euclidean convexity in the following sense. Let $\mathcal{A}=(U,\mathcal{C})$ be a convex geometry. We say that $\mathcal{A}$ is CG-realizable if $U$ can be viewed as a finite subset of points of ${\mathbb{R}}^{d}$ such that $X\subseteq U$ belongs to $\mathcal{C}$ if and only if the Euclidean convex hull of $X$ does not contain other points of $U$: $\operatorname{conv}_{{\mathbb{R}}^{d}}(X)\cap U=X$. Not every convex geometry can be realized in this way. Kashiwabara et al. [19] gave a realization theorem for all convex geometries using the notion of generalized convex shellings. Let $P$ and $Q$ be two finite sets of ${\mathbb{R}}^{d}$ such that the Euclidean convex hull of $Q$ does not intersect $P$. Then the family of subsets $\mathcal{C}(P,Q)=\{X\subseteq P:\operatorname{conv}(X\cup Q)\cap P=X\}$ of $P$ is called the generalized convex shelling on $P$ with respect to $Q$. It is easy to see that any generalized convex shelling is a convex geometry. Conversely, Kashiwabara et al. [19] proved that any convex geometry arises as a generalized convex shelling. If $Q=\varnothing$, then this leads to the notion of CG-realizability. If ${\mathcal{C}}=\mathcal{C}(P,Q)$, then we will say that $\mathcal{C}(P,Q)$ is a generalized convex shelling of $\mathcal{A}=(P,\mathcal{C})$. Richter and Rogers [28] proved that any convex geometry of convex dimension $d$ has a generalized convex shelling in ${\mathbb{R}}^{d}$.
@@
We say that $(P,Q)$ is a point representation of a COM $\mathcal{M}$ if $\mathcal{M}=\mathcal{M}(P,Q)$, i.e., the set $\mathcal{T}$ of topes of $\mathcal{M}$ consists precisely of the $\{-1,+1\}$-vectors of all sets of $\mathcal{M}(P,Q)$ as in Remark 3. Trivially, $P\in\mathcal{M}(P,Q)$. Now, we establish a correspondence between (hyperplane) realizations and point realizations of COMs and a link with generalized convex shellings. For this purpose we define a COM to be acyclic if it has the all plus-vector $(+1,\ldots,+1)$ as a tope.
@@
Let $\mathcal{M}=(U,\mathcal{L})$ be an acyclic realizable COM. We have $\mathcal{M}=\mathcal{M}(\mathcal{H}^{\prime},K^{\prime})$ where ${\mathcal{H}}^{\prime}$ is a central hyperplane arrangement ${\mathcal{H}}^{\prime}$ and $K^{\prime}$ an open polyhedral cone in $\mathbb{R}^{d+1}$ as in Lemma 5. Now, the vectors $v_{e}$, normal to the hyperplanes $H^{\prime}_{e},e\in U$ of the arrangement ${\mathcal{H}}^{\prime}$, gives rise to a set ${\mathcal{V}}$ of vectors of ${\mathbb{R}}^{d+1}$. Analogously, the vectors $u_{j}$, normal to the hyperplanes $S^{\prime}_{j}$ of ${\mathcal{S}}^{\prime}$ defining the polyhedron $K^{\prime}$ and such that their product with the vectors with ends inside $K^{\prime}$ is positive, gives rise to a set ${\mathcal{U}}$ of vectors of ${\mathbb{R}}^{d+1}$. For a cell $C$ of the initial realization, each point $p$ of $C$ defines a vector $v_{p}$ belonging to the cone $C^{\prime}$. Let $H_{v_{p}}$ be the central hyperplane of ${\mathbb{R}}^{d+1}$ having $v_{p}$ as the normal vector. Then all vectors of ${\mathcal{U}}$ belong to a halfspace defined by $H_{v_{p}}$. Analogously, each vector $v_{e}\in{\mathcal{V}}$ belongs to the positive side of the hyperplane $H_{v_{p}}$ if and only if the cell $C$ is located on the positive side of the hyperplane $H_{e}\in{\mathcal{H}}$ (and thus the cone $C^{\prime}$ is on the positive side of the central hyperplane $H^{\prime}_{e}\in{\mathcal{H}}$). Therefore, the topes of $\mathcal{M}$ can be identified (in the set-theoretical language) with the subsets $X$ of $U$ such that there exists a central hyperplane $H^{\prime}$ of ${\mathbb{R}}^{d+1}$ such that all $U$ belong to one halfspace defined by $H^{\prime}$ and a vector $v_{e},e\in U$ belongs to the same halfspace defined by $H^{\prime}$ if and only if $e\in X$.
@@
Therefore, the hyperplane realizations and point representations of COMs are much simpler than generalized convex shellings and correspond to classical OM-realizations. The main results of our paper (Theorem 26 and Theorem 35) must be considered from this point of view, while compared to the results of [19] and [28] in the case of convex geometries. However, in contrast to our results generalized convex shellings characterize convex geometries [19].
In the following we establish a connection between point representations and generalized convex shellings. A COM $\mathcal{M}$ is called intersection-closed if the set of topes of $\mathcal{M}$ defines an intersection-closed family of sets. Note that by simplicity this implies that $(-1,\ldots,-1)$ is a tope of $\mathcal{M}$. Obviously, convex geometries are acyclic and their ideals are intersection closed COMs. Acyclicity corresponds to the universe being convex, i.e., $(+1,\ldots,+1)$ is a tope of $\mathcal{M}$.
Proposition 14.
Let $P,Q\subset\mathbb{R}^{d}$ be finite. If $\mathcal{M}=\mathcal{M}(P,Q)$ is a simple acyclic intersection-closed COM, then $\mathcal{C}(P,Q)$ is a generalized convex shelling. In particular, $\mathcal{M}$ is a convex geometry.
@@
Since $\mathcal{M}=\mathcal{M}(P,Q)$ contains the tope $(-1,\ldots,-1)$, necessarily $\operatorname{conv}(Q)\cap P=\varnothing$. From the definition of the point representations and generalized convex shellings, we conclude that $\mathcal{M}\subseteq\mathcal{C}(P,Q)$. To prove the converse inclusion $\mathcal{C}(P,Q)\subseteq\mathcal{M}$, pick any set $X$ from $\mathcal{C}(P,Q)$. Then there exists a set $H^{+}_{1},\ldots,H^{+}_{k}$ of halfspaces, all containing $Q$ and such that $X=P\cap(\cap H_{i}^{+})$. Let $X_{i}=P\cap H_{i}^{+},i=1,\ldots,k$. Then each $X_{i}$ belongs to $\mathcal{M}(P,Q)$. Since $\mathcal{M}(P,Q)=\mathcal{M}$, each $X_{i}$ belongs to $\mathcal{M}$. Since $\mathcal{M}$ is intersection-closed and $\cap_{i=1}^{k}X_{i}=X$, we conclude that $X\in\mathcal{M}$, establishing that $\mathcal{C}(P,Q)\subseteq\mathcal{M}$. ∎
@@ 4.1 Rooted circuits
Korte and Lovász [23] (see also the book [24]) identified a canonical subset of rooted circuits, which they call critical rooted circuits. In terms of convex geometries, a rooted circuit $(C,a)$ of a convex geometry $\mathcal{A}=(U,\mathcal{C})$ is a critical rooted circuit if $\operatorname{conv}(C)\setminus\{a\}$ does not belong to $\mathcal{C}$ but $\operatorname{conv}(C)\setminus\{a,b\}$ belongs to $\mathcal{C}$ for any $b\in C\setminus\{a\}$. Denote by $\mathcal{R}_{0}(\mathcal{A})$ the set of all critical rooted circuits of $\mathcal{A}$. The importance of the family $\mathcal{R}_{0}(\mathcal{A})$ of critical rooted circuits stems from the fact that $\mathcal{A}$ can be uniquely determined by $\mathcal{R}_{0}(\mathcal{A})$ and $\mathcal{R}_{0}(\mathcal{A})$ is minimal with respect to this property. This is: if $\mathcal{A}$ can be determined by $\mathcal{R}^{\prime}\subseteq\mathcal{R}(\mathcal{A})$, then $\mathcal{R}_{0}(\mathcal{A})\subseteq\mathcal{R}^{\prime}$ [23, 24] (for a precise definition of “determined by”, see [24]).
@@ 4.2 Positive circuits
Let $\mathcal{C}^{\prime\prime}=\{X\in\mathcal{C}:\forall P\in\mathcal{P}(\mathcal{I}),P\not\subseteq X\}$. Pick any $X\in\mathcal{C}^{\prime}$ and any $P\subseteq X$. Then $P\cap X=P\in\mathcal{C}^{\prime}_{|P}$ and thus $\mathcal{C}^{\prime}_{|P}\neq 2^{P}\setminus\{P\}$. This establishes that $P\notin\mathcal{P}(\mathcal{I})$ for any $P\subseteq X$ and consequently $X\in\mathcal{C}^{\prime\prime}$. Therefore $\mathcal{C}^{\prime}\subseteq\mathcal{C}^{\prime\prime}$.
Suppose now that there exists $Y\in\mathcal{C}^{\prime\prime}\setminus\mathcal{C}^{\prime}$ and assume that $Y$ is such a set of minimal size. By minimality of $Y$, for any $e\in Y$, $Y\setminus\{e\}\in\mathcal{C}^{\prime\prime}$ if and only if $Y\setminus\{e\}\in\mathcal{C}^{\prime}$. Let $P_{0}=\{e\in Y:Y\setminus\{e\}\in\mathcal{C}^{\prime}\}$. Since $\mathcal{C}^{\prime}$ is intersection-closed, for any $\varnothing\subsetneq P^{\prime}\subseteq P_{0}$, we have $Y\setminus P^{\prime}\in\mathcal{C}^{\prime}$. This shows that $2^{P_{0}}\setminus\{P_{0}\}\subseteq\mathcal{C}^{\prime}_{|P_{0}}\subseteq 2^{P_{0}}$. Since $P_{0}\subseteq Y$ and since $Y\in\mathcal{C}^{\prime\prime}$, by the definition of $\mathcal{C}^{\prime\prime}$, we get $P_{0}\notin\mathcal{P}(\mathcal{I})$ and thus $\mathcal{C}^{\prime}_{|P_{0}}=2^{P_{0}}$ by the definition of $\mathcal{P}(\mathcal{I})$. Consequently, there exists $X\in\mathcal{C}^{\prime}$ such that $P_{0}\subseteq X$. Let $Z=X\cap Y$ and observe that $P_{0}\subseteq Z$. Since $\mathcal{C}$ is intersection-closed, $Z\in\mathcal{C}$. If $Z=Y$, then $Y\subseteq X\in\mathcal{C}^{\prime}$, contradicting the fact that $\mathcal{C}^{\prime}$ is an ideal of $\mathcal{C}$. Thus we have $Z\subsetneq Y$. Since $Z\in\mathcal{C}^{\prime}$ and $Y\notin\mathcal{C}^{\prime}$, there exists a set $Z\subseteq A\subseteq Y$ and an element $e\in Y\setminus A$ such that $A\in\mathcal{C}^{\prime}$ and $B=A\cup\{e\}\in\mathcal{C}\setminus\mathcal{C}^{\prime}$ ($A$ and $e$ exist since $\mathcal{C}$ is ample and thus isometric). By minimality of $Y$, $B=Y$. Consequently $e\in P_{0}$ by the definition of $P_{0}$, but this is impossible since $P_{0}\subseteq Z\subseteq A$. This establishes that $\mathcal{C}^{\prime\prime}\subseteq\mathcal{C}^{\prime}$. ∎